OpenStudy (anonymous):

Determine the parametric equations of the position of a particle with constant velocity that follows a straight line path on the plane if it starts at the point P (9,7) and after one second it is at the point Q(−7,10) . I need to find x(t), y(t) and the speed of the particle. I know that vector PQ = <-16,3> Would my parametric eqs be x(t) = 9 -16t, and y(t) = 7 + 3t ???

OpenStudy (anonymous):

P=<9,7> Q=<-7,10>

OpenStudy (anonymous):

Q-P

OpenStudy (anonymous):

Q-P = <-16,3> right?

OpenStudy (anonymous):

Please someone provide a clear explanation

OpenStudy (anonymous):

Yes you are right

OpenStudy (anonymous):

Now p is your starting point,right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

<-16,3> is the direction

OpenStudy (anonymous):

so my equations posted above are correct?

OpenStudy (anonymous):

wouldn't it be 7 +3t?

OpenStudy (anonymous):

So we add origin to direction 9-16t 7+3t

OpenStudy (anonymous):

Can anyone explain this to me verbally with detail?

OpenStudy (anonymous):

So if you plug in 1 for t you would get Q

OpenStudy (anonymous):

So your are starting at P and moving in the direction of Q

OpenStudy (anonymous):

I understand that, I am asking if my equations posted in the original post are correct, and if so, how do I determine the speed. I know to use t=1, and plug that in to both of my equations, but after that I am confused. Once I get the values for x(t) and y(t) I am unsure of how to find the speed.

OpenStudy (anonymous):

Your equation is correct. To Find speed you will have to take a derivative x(t) = 9 -16t, and y(t) = 7 + 3t ??? x[t]'=-16 y[t]'=3